2. (12 pts.) Find the exact value. a. sin(165) c. sin(52.5°) sin(7.5°) b. tan(15) 1 -tan (15)
Answer #1

@ usin (165) = sin (120°+45°) > sin (120).cos(459) + cos204) sin(45) sin (120°) = sin(90* +3) cos(120) = cool goº +30°) = c

sin (52-5). Sin (7.5) = sin (459 +7.5), sin (7:5) =[smiles).cos(7-9) +003/1.59.sam(7.5)] sin(7:5) [ta cas (7.5) + E si1, 2 ศ (35)(35) - 22 sinfto sin (7.5). cos 8(1-s) = 13 - ) H2 but in eq. 0 ป3 - 1 12 + 22 - 3 -| หป ป2 11 ป) 22 - 2 หาป2 11 2